OpenStudy (anonymous):

need some more help... A (-10, -2) and B(2,-8) are the endpoints of a diameter of a circle. Find the radius of the circle.

jimthompson5910 (jim_thompson5910):

radius = half of diameter in this case, diameter = distance from A to B So we need the distance from A to B. To do that, we use the distance formula d = sqrt((x2-x1)^2+(y2-y1)^2) Remember that the two generic points are (x1, y1) and (x2, y2) So (x1, y1) = (-10, -2), which means x1=-10 and y1= -2 and (x2, y2) = (2, -8), which means x2=2 and y2=-8 So in short, x1=-10, y1= -2, x2=2, and y2=-8. ------------------- d = sqrt((x2-x1)^2+(y2-y1)^2) d = sqrt((2-(-10))^2+(-8-( -2))^2) d = sqrt((2+10)^2+(-8-( -2))^2) d = sqrt((12)^2+(-6)^2) d = sqrt(144+36) d = sqrt(180) d = 6*sqrt(5) So exact distance between the two points is 6*sqrt(5) units. So the diameter is 6*sqrt(5) units Cut this value in half to get 3*sqrt(5) units So the radius is 3*sqrt(5) units

OpenStudy (anonymous):

that's not one of the answers. the answers are: (a) 6 radical 5 over 2, (b) 2 radical 41, (c) 6i, (d) -3, (e) 6 radical 5

jimthompson5910 (jim_thompson5910):

well then it's choice A since they didn't reduce (not sure why?)

jimthompson5910 (jim_thompson5910):

but 6 radical 5 over 2 is the same as 3*sqrt(5) which is what I got

OpenStudy (anonymous):

thanks a bunch man